OpenStudy (anonymous):

four less than a quotient of a number and 6 is 3. Find the number

OpenStudy (anonymous):

42?

OpenStudy (anonymous):

\[4 -{x \over 6} = 3\]

OpenStudy (anonymous):

-x/6 = 3-4 -x/6 = -1 x = 6

OpenStudy (anonymous):

thanks